8 (x+1) + 2 (3x+1)=160


1
Expert's answer
2021-04-14T07:28:08-0400

Step 1

Open the brackets

8x+8+6x+2=160

Step 2

collect like terms and perform mathematical operation

8x+6x+8+2=160

14x+10=160

Step 3

Subtract 10 from both sides

14x+10-10=160-10

14x=150

Step 4

Divide both sides by 14

"\frac {14x}{14}=\frac{150}{14}"

x="\frac {75}{7}=10\frac 5 7"